OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of and passes through the point (2, -7).

OpenStudy (ivettef365):

what is the slope ?

OpenStudy (anonymous):

2/3

OpenStudy (ivettef365):

use the formula y = mx + b to find b you have (x,y) = (2,-7) you have m = slope = 2/3

OpenStudy (anonymous):

once i find b, what do i do?

OpenStudy (ivettef365):

then you have the equation of the line in slope-form all you need is the m and the b, the x,y stays as it is, because x,y could be any points on that line

OpenStudy (ivettef365):

can you find the b, and I show you how to write the equation

OpenStudy (anonymous):

8.33 or 8 and 1/3

OpenStudy (anonymous):

2/3

OpenStudy (ivettef365):

how did you get 2/3 ?

OpenStudy (ivettef365):

y = mx + b -7 = (2/3)(2) + b

OpenStudy (ivettef365):

-7 = 4/3 + b b = -7 -(4/3) b = 7 4 -7(3) -4 -25 - --- - --- = --------- = ------- 1 3 3 3

OpenStudy (ivettef365):

so you have b = -25/3 and m = 2/3 now you can write the equation of the line as follows y = 2/3 x -25/3